is 3d 7 more stable than 4s 1 ?if yes,reason???

is 3d7 more stable than 4s1 ?if yes,reason???

Grade:12

1 Answers

Aarti Gupta
9 years ago
4s1 will be more stable than 3d7.This can be understand by Bohr- Burry rule,which is also known as (n+l) rule.According to this rule-
A sub-shell with lower value of (n+l) has lower energy.
Thus,
in 4s orbital n = 4 and l = 0 ,hence n+ l = 4+ 0 = 4
in 3d orbital n = 3 and l = 2 ,hence, n + l = 3 + 2 = 5
so,4s orbital has lower energy than 3d-orbital and we know that lower energy means more stability.Thus 4s orbital having lower energy will be more stable than 3d orbital with more energy.